Visit your local library: On a recent Saturday, a total of 1480 people visited a local library. Of these people, 251were under age10, 453 were aged 10-18, 165 were aged 19-30, and the rest were more than 30 years old. One person is sampled at random.

(a) What is the probability that the person is less than 19 years old?Round your answer to four decimal places.

(b) What is the probability that the person is more than 30 years old? Round your answer to four decimal places.
Answer #1

answer)

Probability is given by favorable/total

Total here is 1480

A)

Favorable = less than 19 years old = 453 + 251 = 704

Required probability is 704/1480

= 0.47567567567

= 0.4757

B)

Favorable = more than 30 years old = 1480 - (251+453+165) = 611

Required probability is 611/1480

= 0.41283783783

= 0.4128